BestDeal Inc. has just paid dividend of $5.The dividend is expected to increase at 16% for the

first 15 years, -6% for the next 10 years and then 2% forever. The required rate of return is 12% for the first 8 years and 20% thereafter. Find its share price today.
Answer #1
Share Price today=present Value of Future Cash Flows
Present Value (PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=Discount Rate=Required rate of return
N=Year of Cash Flow
D0 Current Dividend $5
D1 Dividend in Year1=D0*(1+0.16)=5*1.16 $5.80
Dn+1 Dividend in Year(n+1)=Dn*(1+0.16)
(For n< or =14
D16=D15*(1-0.06)=D15*0.94
For n>or=16 and n< or=24
Dn+1=Dn*0.94
D26=D25*(1+0.02)
For n>or=25
Dn+1=Dn*1.02
Year wise Dividend is given below:
Year Dividend
D1 1 $5.80
D2 2 $6.73
D3 3 $7.80
D4 4 $9.05
D5 5 $10.50
D6 6 $12.18
D7 7 $14.13
D8 8 $16.39
D9 9 $19.01
D10 10 $22.06
D11 11 $25.59
D12 12 $29.68
D13 13 $34.43
D14 14 $39.94
D15 15 $46.33
D16 16 $43.55
D17 17 $40.94
D18 18 $38.48
D19 19 $36.17
D20 20 $34.00
D21 21 $31.96
D22 22 $30.04
D23 23 $28.24
D24 24 $26.55
D25 25 $24.95
D26 26 $25.45
Required rate of Return in year 26 onwards 0.2 20%
Dividend Growth rate from year 26 onwards 0.02 2%
P25 Price at end of Year 25=D26/(0.2-0.02)=25.45/0.18 $141.40
Required return from year 9 onwards 20% 0.2
Let us Calculated Present Value of cash flows at end of year 8
Consider Year 8 as 0year and year 9 asyear1 and so on
N CF PV=CF/(1.2^N)
Year Cash Flow Present value at end of year8
D9 1 $19.01 15.84567198
D10 2 $22.06 15.31748291
D11 3 $25.59 14.80690015
D12 4 $29.68 14.31333681
D13 5 $34.43 13.83622558
D14 6 $39.94 13.37501806
D15 7 $46.33 12.92918413
D16 8 $43.55 10.1278609
D17 9 $40.94 7.933491039
D18 10 $38.48 6.21456798
D19 11 $36.17 4.868078251
D20 12 $34.00 3.813327963
D21 13 $31.96 2.987106905
D22 14 $30.04 2.339900409
D23 15 $28.24 1.832921987
D24 16 $26.55 1.43578889
D25 17 $24.95 1.124701297
P25 17 $141.40 6.373307349
SUM $149.47
Present value in Year 8=Price in year8=P8 $149.47
Required return in Year 1 to8 12% 0.12
PRESENT VALUE AT END OF YEAR 0
N CF PV=CF/(1.12^N)
Year Cash Flow Present value at end of year0
D1 1 $5.80 5.178571429
D2 2 $6.73 5.363520408
D3 3 $7.80 5.555074708
D4 4 $9.05 5.753470234
D5 5 $10.50 5.958951314
D6 6 $12.18 6.171771003
D7 7 $14.13 6.392191396
D8 8 $16.39 6.620483946
P8 8 $149.47 60.37039404
SUM $107.36
SHARE PRICE TODAY $107.36